# Ultra-low-power ring-based wireless mouse

### v1

@@ 4.2. SNR of picoRing mouse
We first evaluate the signal-to-noise ratio (SNR) of picoRing mouse for user’s scrolling and pressing interactions, input power, distance between the ring and wristband, the finger bending, and public environments. Similar to (Takahashi et al., 2024), the SNR was calculated as follows using S21 outputs of the VNA:
Here, we consider the SNR over can support high-fidelity recognition over accuracy, referring to (Takahashi et al., 2024). Also, we decouple the readout board with the common ground using the RF transformers (T1-1-X65+, Mini Circuits). Figure 6a shows the time-series SNR and the average SNR for the pressing and 2D scrolling inputs. The average SNR was calculated using data collected over a 5-minute period from three invited participants with almost the similar hand sizes (one woman and two men, 20s). The result shows that the average SNR for each input is over with minimal variations, indicating that picoRing mouse can support high-fidelity recognition of thumb-to-mouse finger inputs for the similar hand size. To be available for various hand sizes, we will quantitatively explore how the coil size can affect the SNR using those like electromagnetic simulators.
Next, we measured how the distance and orientation of the ring coil influence the SNR of picoRing mouse. picoRing (Takahashi et al., 2024) shows almost the SNR characteristics with or without the hands, thanks to the robustness of the ring-to-wristband inductive coupling against the dielectric body, we conducted this evaluation without human hands to easily adjust the distance and orientation between the coils. Specifically, we prepared a set of jig to change the distance and orientation of the ring coil by the 3D printer (see Figure 6b). The range of he distance is from to in step of in addition to varying the angle of the ring coil from to in step of . We prepared the two types of the ring coils: straight-type coil and tilted-type coil with the coil angled about to the vertical axis of the ring. Because the thumb-to-index input, which is typically performed with the finger in the bent position, can cause misalignment between the ring and wristband coils, weakening magnetic coupling, the tilted ring with the coil angled to the vertical axis enables to maintain the alignment and strengthen the inductive coupling. Note that the ring’s resonant frequency is tuned at . The result shows that the available distance is estimated to be approximately because the SNR is over against up to finger bending. These available distance and angle provided by picoRing mouse are well-suited for the reliable thumb-to-index input during the typical middle-sized hand movements. Furthermore, the tilted ring could increase the SNR by approximately when the finger is bent by around because the measured inductive coupling coefficient () increases from to .
@@
Finally, we measured the SNR during the daily usage. Although the SNRs for proximity to metallic items and electrical appliances are shown to be robust in picoRing (Takahashi et al., 2024), our focus is on measuring the SNR in the real-world wearable computing scenarios. Here, we consider three practical scenarios: using the ring while 1) seated in a train or airplane, 2) standing in crowded indoors or outdoors, and 3) driving a car or riding a bicycle. Figure 7c shows the time-series SNR spectrum and average SNR for these situations, based on data from a single user (20s, man). The result indicates that picoRing mouse successfully detects a sharp peak at despite proximity to the metallic items (e.g., steering wheels) and electrical appliances (e.g., laptops, smartphones, electronic control units in the cars and trains).
@@ 5. APPLICATION EXAMPLES
picoRing mouse offers a significant advantage in providing long-term continuous operation on a single charge, while reliably detecting subtle thumb-to-index inputs. Therefore, its primary application is ubiquitous finger-based interaction systems, particularly when integrated with augmented reality (AR) glasses or head-mounted displays (HMD) (see Figure 8). AR glasses and HMDs often rely on dynamic hand gestures for scrolling interactions, which can lead to fatigue over time. In contrast, picoRing mouse offers subtle finger inputs that enable comfortable, long-term interactions while preserving user privacy. This makes it an ideal solution for seamless engagement with digital content. Additionally, the device’s efficient power management allows frequent and natural use without concerns about battery life, enhancing its practicality in everyday AR and HMD applications. Figure 8 shows some application examples of picoRing mouse. The discreet nature of ring inputs are not easily noticed by surroundings, minimizing any disruption to those nearby. For example, in an airplane or train, a user discreetly interacts with a personal screen using the picoRing mouse, allowing for private control without disturbing nearby passengers. Even in a urban crowded space, users navigate a large public display or AR map with the picoRing mouse, seamlessly interacting with digital content.
@@ 6. DISCUSSION
Dual-device requirement. picoRing mouse requires users to attach the two types of wearable devices—ring and wristband—unlike the standard wearables like smartwatches and smartrings. Given that many users already wear wristbands throughout the day for health and fitness tracking, as well as for convenient visual displays, the burden of wearing an additional ultra-low-power ring that requires minimal charging is quite light, making it easy to incorporate the wearing a pair of the ring and the wristband into daily routines.
@@
Bulky and power-consuming readout board. Currently, our readout board, which includes the bridge, VNA, and the external PC, requires a smartphone-sized NanoVNA. However, the board could be miniaturized to fit into a smartwatch. By utilizing chip-sized VNAs (ADL5960, Analog Devices) and high-performance microcontrollers, we can significantly reduce the size to as small as and power consumption to as low as while maintaining functionality. Instead of the VNA, the use of the customized NFC protocols (Takahashi et al., 2025), activating only the listening mode in NFC protocol to capture the frequency response from the NFC sensor tag without wireless power supply, could decrease the total power consumption less than (KG, [n. d.]; Zhao et al., 2015). Such a low power operation of the wristband coil could allow the integration of our wristband system into the commercially-available smartwatch, which is typically driven by about -class Lipo battery (González-Cañete and Casilari, 2021).
@@
Wireless charging while in use. Wearable devices often recharge naturally when not in use, such as earbuds placed in dedicated charging cases. However, rings and wristbands are typically worn continuously, making this more challenging. The integration of wearable power transmission system such as charging-enabled textiles could achieve the continuous power supply to picoRing mouse in the use (Sato et al., 2025; Takahashi et al., 2022; Li et al., 2025b). This would ensure that picoRing mouse remains powered throughout daily activities, enabling to be worn continuously as a fashion accessory.
@@
Limited user study. In the current study, we primarily focused on measuring the SNR related to mouse input. Future investigations will incorporate a multi-factorial repeated-measures ANOVA to assess additional aspects such as ease of use, social acceptance, and fatigue (Chen et al., 2022). We plan to conduct a comprehensive evaluation of subjective usability, utilizing tools such as the System Usability Scale (SUS) and the NASA Task Load Index (NASA-TLX). These evaluations will provide further understanding of the picoRing mouse in various real-world scenarios, helping to refine its design and improve user experience.
@@
Extension of ring functionality. picoRing mouse has the potential to become a versatile ultra-low-powered wearable platform for a wide range of applications by integrating various low-powered sensors (Wang et al., 2025). For instance, by incorporating IMUs (Shen et al., 2024; Zhao et al., 2022) or speaker (Yu et al., 2024), picoRing can be transformed into motion sensing rings (see Figure 9). Additionally, integrating PPG or temperature sensors could transform it into a healthcare rings. Furthermore, wearing picoRing mouse on the both hands can enable wearable VR/AR interface even in the public spaces, by converting the user’s subtle mouse input to extended 3D hand dynamic posture (Kari and Holz, 2023).
Cross-device interactions by ring and wristband. picoRing system is limited to input from the ring alone. This limitation could be solved by installing the wristband-based hand sensing system (Dementyev and Paradiso, 2014; Kim and Harrison, 2022) into the wristband coil. Such a hybrid system can enable the recognition of both subtle finger inputs and dynamic hand gestures.
@@
Optimization for further ring’s low-powered operation. We are utilizing the ultra-low-powered MCU (STM32L011F4U6, STM), and there is still room for reducing power consumption in both hardware and software aspects. For instance, we can lower the operating voltage from to or reduce the MCU’s clock speed. We would explore these options to increase further battery runtime, resulting in the decrease of the battery size.
@@
Error handling in frequency-shift keying. The current frequency-shift keying in our ring-to-wristband communication is so simple, using different frequencies to send the digital information from the mouse sensor. This simplicity presents several challenges such as security concerns and data integrity issues. Furthermore, the slow data rate of bps in current picoRing mouse allows the high SNR against electromagnetic noise from nearby devices. However, the fast communication at kbps-level such as for IMU data would make picoRing mouse vulnerable to electromagnetic noise, causing a significant drop in SNR. To ensure reliable and fast ring-to-wristband communication, error handling techniques such as forward error correction would be commonly available.
@@ 7. CONCLUSION
This paper presents picoRing mouse, an ultra-low-power ring-based finger input device. By combining semi-PIT-based ring-to-wristband low-powered wireless communication with the low-powered mouse module, picoRing mouse achieves long-term operation of approximately 600 (8hrs use/day)-1000 (4hrs use/day) hours, in addition to supporting reliable, multi-modal, subtle inputs. Our technical evaluation shows the strong robustness of ring-to-wristband wireless link against different users, hand postures, and potential situations in the field of wearable computing. We strongly believe picoRing mouse could provide a new class of long-term wearables, seamlessly interconnecting itself with daily HCI devices.
@@ Acknowledgements.
This work was mainly supported by JST JPMJAX21K9, JSPS KAKEN 22K21343, JST ASPIRE JPMJAP2401, and the Asahi Glass Foundation.References
@@
- Gummeson et al. (2014) Jeremy Gummeson, Bodhi Priyantha, and Jie Liu. 2014. An energy harvesting wearable ring platform for gestureinput on surfaces. In Proceedings of the 12th annual international conference on Mobile systems, applications, and services (MobiSys ’14). Association for Computing Machinery, New York, NY, USA, 162–175. https://doi.org/10.1145/2594368.2594389
@@
- Ogata et al. (2012) Masa Ogata, Yuta Sugiura, Hirotaka Osawa, and Michita Imai. 2012. iRing: intelligent ring using infrared reflection. In Proceedings of the 25th annual ACM symposium on User interface software and technology (UIST ’12). Association for Computing Machinery, New York, NY, USA, 131–136. https://doi.org/10.1145/2380116.2380135
@@
- Streli et al. (2024) Paul Streli, Mark Richardson, Fadi Botros, Shugao Ma, Robert Wang, and Christian Holz. 2024. TouchInsight: Uncertainty-aware Rapid Touch and Text Input for Mixed Reality from Egocentric Vision. In Proceedings of the 37th Annual ACM Symposium on User Interface Software and Technology. ACM, Pittsburgh PA USA, 1–16. https://doi.org/10.1145/3654777.3676330
@@
- Waghmare et al. (2023) Anandghan Waghmare, Youssef Ben Taleb, Ishan Chatterjee, Arjun Narendra, and Shwetak Patel. 2023. Z-Ring: Single-Point Bio-Impedance Sensing for Gesture, Touch, Object and User Recognition. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (CHI ’23). Association for Computing Machinery, New York, NY, USA, 1–18. https://doi.org/10.1145/3544548.3581422
@@
- Wang et al. (2025) Zeyu Wang, Ruotong Yu, Xiangyang Wang, Jiexin Ding, Jiankai Tang, Jun Fang, Zhe He, Zhuojun Li, Tobias Röddiger, Weiye Xu, Xiyuxing Zhang, huan-ang Gao, Nan Gao, Chun Yu, Yuanchun Shi, and Yuntao Wang. 2025. Computing with Smart Rings: A Systematic Literature Review. https://doi.org/10.48550/arXiv.2502.02459 arXiv:2502.02459 [cs].
@@
- Zhou et al. (2023) Hao Zhou, Taiting Lu, Yilin Liu, Shijia Zhang, Runze Liu, and Mahanth Gowda. 2023. One Ring to Rule Them All: An Open Source Smartring Platform for Finger Motion Analytics and Healthcare Applications. In Proceedings of the 8th ACM/IEEE Conference on Internet of Things Design and Implementation (IoTDI ’23). Association for Computing Machinery, New York, NY, USA, 27–38. https://doi.org/10.1145/3576842.3582382

### v2

@@ 4.2. SNR of picoRing mouse
We then evaluate the signal-to-noise ratio (SNR) of picoRing mouse for user’s scrolling and pressing interactions, input power, distance between the ring and wristband, the finger bending, and public environments. Similar to (Takahashi et al., 2024), the SNR was calculated as follows using S21 outputs of the VNA:
Here, we consider that the SNR over can support high-fidelity recognition over accuracy, referring to (Takahashi et al., 2024). Also, we decouple the readout board with the common ground using the RF transformers (T1-1-X65+, Mini Circuits). Figure 6a shows the time-series SNR and the average SNR for the pressing and 2D scrolling inputs. The average SNR was calculated using data collected over a 5-minute period from three invited participants with almost similar hand sizes (one woman and two men, 20s). The result shows that the average SNR for each input is over with minimal variations, indicating that picoRing mouse can support high-fidelity recognition of thumb-to-mouse finger inputs for a similar hand size. To be available for various hand sizes, we will quantitatively explore how the coil size can affect the SNR using those like electromagnetic simulators.
Next, we measured how the distance and orientation of the ring coil influence the SNR of picoRing mouse. Original picoRing (Takahashi et al., 2024) have shown that the SNR characteristics is almost the same with or without the hands, thanks to the inductive inherent robustness to dielectric human presence. Therefore, we conducted this evaluation without human hands to easily adjust the distance and orientation between the coils. Specifically, we prepared a set of jigs to change the distance and orientation of the ring coil by the 3D printer (see Figure 6b). The range of he distance is from to in step of in addition to varying the angle of the ring coil from to in step of . We prepared the two types of ring coils: straight-type coil and tilted-type coil with the coil angled about to the vertical axis of the ring. Because the thumb-to-index input, which is typically performed with the finger in the bent position, can cause misalignment between the ring and wristband coils, weakening magnetic coupling, the tilted ring with the coil angled to the vertical axis enables to maintain the alignment and strengthen the inductive coupling. Note that the ring’s resonant frequency is tuned at . The result shows that the available distance is estimated to be approximately because the SNR is over against up to finger bending. These available distance and angle provided by picoRing mouse are well-suited for the reliable thumb-to-index input during the typical middle-sized hand movements. Furthermore, the tilted ring could increase the SNR by approximately when the finger is bent by around because the measured inductive coupling coefficient () increases from to .
@@
Finally, we measured the SNR during the daily usage. Although the SNRs for proximity to metallic items and electrical appliances are shown to be robust in picoRing (Takahashi et al., 2024), our focus is on measuring the SNR against electromagnetically-noisy situations in real-world wearable computing scenarios. Here, we consider the following three situations: using the ring while 1) seated in a train, 2) standing in crowded areas in addition to being surrounded by the multiple smartphones, and 3) driving a car. Note that the measured noise level in these situations is approximately higher than in typical environments. Figure 7 shows the time-series SNR spectrum and average SNR for these situations, based on data from a single user (20s, man). The result indicates that picoRing mouse successfully detects a sharp peak at with the sufficient SNR over despite proximity to the metallic items (e.g., steering wheels) and electrical appliances (e.g., laptops, smartphones, electronic control units in the cars and trains).
@@ 5. APPLICATION EXAMPLES
picoRing mouse offers a significant advantage in providing long-term continuous operation on a single charge, while reliably detecting subtle thumb-to-index inputs. Therefore, its primary application is ubiquitous finger-based interaction systems, particularly when integrated with augmented reality (AR) glasses or head-mounted displays (HMD) (see Figure 8). AR glasses and HMDs often rely on dynamic hand gestures for scrolling interactions, which can lead to fatigue over time. In contrast, picoRing mouse offers subtle finger inputs that enable comfortable, long-term interactions while preserving user privacy. This makes it an ideal solution for seamless engagement with digital content. Additionally, the device’s efficient power management allows frequent and natural use without concerns about battery life, enhancing its practicality in everyday AR and HMD applications. Figure 8 shows some application examples of picoRing mouse. The discreet nature of ring inputs are not easily noticed by the surroundings, minimizing any disruption to those nearby. For example, in an airplane or train, a user discreetly interacts with a personal screen using the picoRing mouse, allowing for private control without disturbing nearby passengers. Even in an urban crowded space, users navigate a large public display or AR map with the picoRing mouse, seamlessly interacting with digital content (https://youtu.be/O7E_RQ8KSLQ).
This figure presents ring-sized sensors for input, motion, and healthcare, categorized by power consumption: battery-free, uW-class, and mW-class. For input and motion, battery-free sensors include a mechanical switch for press and slide actions and a magnet for motion detection. The uW-class features a 2D trackball/joystick for scrolling and a 3 or 6-axis IMU for motion sensing. The mW-class includes a camera and microphone for motion detection, along with a 9-axis IMU. For healthcare, the uW-class has a temperature sensor for stress level monitoring, while the mW-class includes a PPG sensor for circulatory status assessment. The layout highlights the power requirements and functionalities of each sensor type.
@@ 6. DISCUSSION
Dual-device requirement. picoRing mouse requires users to attach the two types of wearable devices—ring and wristband—unlike the standard wearables like smartwatches and smart rings. Given that many users already wear wristbands throughout the day for health and fitness tracking, as well as for convenient visual displays, the burden of wearing an additional ultra-low-power ring that requires minimal charging is quite light, making it easy to incorporate wearing a pair of the ring and the wristband into daily routines.
@@
Bulky and power-consuming readout board. Currently, our readout board, which includes the bridge, VNA, and the external PC, requires a smartphone-sized NanoVNA. However, the board could be miniaturized to fit into a smartwatch. NFC, which hardware is optimized for the coil-based communication (Takahashi et al., 2025), could decrease the total power consumption less than (KG, [n. d.]; Zhao et al., 2015). Such a low power operation of the wristband coil could allow the integration of our wristband system into the commercially-available smartwatch, which is typically driven by about -class Lipo battery (González-Cañete and Casilari, 2021).
@@
Wireless charging while in use. Wearable devices often recharge naturally when not in use, such as earbuds placed in dedicated charging cases. However, rings and wristbands are typically worn continuously, making this more challenging. The integration of a wearable power transmission system such as charging-enabled textiles could achieve the continuous power supply to picoRing mouse in the use (Sato et al., 2025; Takahashi et al., 2022; Li et al., 2025b). This would ensure that picoRing mouse remains powered throughout daily activities, enabling to be worn continuously as a fashion accessory.
@@
Limited user study. In the current study, we primarily focused on measuring the SNR related to mouse input. Unlike antenna-based wearable devices operating at higher frequencies (e.g., 2.4 GHz or 920 MHz) which are significantly affected by the human body’s high dielectric loss, our coil-based system experiences almost no electromagnetic interference from a person. This means the signal-to-noise ratio (SNR) and frequency response remain largely unchanged whether the device is worn or not by human hand. One potential user study will be a multi-factorial repeated-measures ANOVA to assess additional aspects such as ease of use, social acceptance, and fatigue (Chen et al., 2022). These evaluations will provide further understanding of the picoRing mouse in various real-world scenarios, helping to refine its design and improve user experience.
@@
Extension of ring functionality. picoRing mouse has the potential to become a versatile ultra-low-powered wearable platform for a wide range of applications by integrating various low-powered sensors. For instance, by incorporating IMUs (Shen et al., 2024; Zhao et al., 2022) or speaker (Yu et al., 2024), picoRing can be transformed into motion sensing rings (see Figure 9). Additionally, integrating PPG or temperature sensors could transform it into a healthcare ring. Furthermore, wearing picoRing mouse on both hands can enable wearable VR/AR interface even in public spaces, by converting the user’s subtle mouse input to extended 3D hand dynamic posture (Kari and Holz, 2023).
@@
Optimization for further ring’s low-powered operation. We are utilizing the ultra-low-powered MCU (STM32L011F4U6, STM), and there is still room for reducing power consumption in both hardware and software aspects. For instance, we can lower the operating voltage from to or reduce the MCU’s clock speed. We would explore these options to increase further battery runtime, resulting in a decrease of the battery size.
@@
Error handling in frequency-shift keying. The current frequency-shift keying in our ring-to-wristband communication is so simple, using different frequencies to send the digital information from the mouse sensor. This simplicity presents several challenges, such as security concerns and data integrity issues. Furthermore, the slow data rate of bps in the current picoRing mouse allows the high SNR against electromagnetic noise from nearby devices. However, the fast communication at kbps-level such as for IMU data would make picoRing mouse vulnerable to electromagnetic noise, causing a significant drop in SNR. To ensure reliable and fast ring-to-wristband communication, error handling techniques such as forward error correction would be commonly available.
@@ 7. CONCLUSION
This paper presents picoRing mouse, an ultra-low-power ring-based finger input device. By combining semi-PIT-based ring-to-wristband low-powered wireless communication with the low-powered mouse module, picoRing mouse achieves long-term operation of approximately 600 (8hrs use/day)-1000 (4hrs use/day) hours, in addition to supporting reliable, multi-modal, subtle inputs. Our technical evaluation shows the strong robustness of the ring-to-wristband wireless link against different users, hand postures, and potential situations in the field of wearable computing. We strongly believe picoRing mouse could provide a new class of long-term wearables, seamlessly interconnecting itself with daily HCI devices.
@@ Acknowledgements.
This work was supported by JST ACT-X JPMJAX21K9, JSPS KAKEN 22K21343, JST ASPIRE JPMJAP2401, and the Asahi Glass Foundation.References
@@
- Gummeson et al. (2014) Jeremy Gummeson, Bodhi Priyantha, and Jie Liu. 2014. An energy harvesting wearable ring platform for gestureinput on surfaces. In Proceedings of the 12th annual international conference on Mobile systems, applications, and services. ACM, Bretton Woods New Hampshire USA, 162–175. https://doi.org/10.1145/2594368.2594389
@@
- Waghmare et al. (2023) Anandghan Waghmare, Youssef Ben Taleb, Ishan Chatterjee, Arjun Narendra, and Shwetak Patel. 2023. Z-Ring: Single-Point Bio-Impedance Sensing for Gesture, Touch, Object and User Recognition. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems. ACM, Hamburg Germany, 1–18. https://doi.org/10.1145/3544548.3581422
@@
- Zhou et al. (2023) Hao Zhou, Taiting Lu, Yilin Liu, Shijia Zhang, Runze Liu, and Mahanth Gowda. 2023. One Ring to Rule Them All: An Open Source Smartring Platform for Finger Motion Analytics and Healthcare Applications. In Proceedings of the 8th ACM/IEEE Conference on Internet of Things Design and Implementation. ACM, San Antonio TX USA, 27–38. https://doi.org/10.1145/3576842.3582382